CM 体の Hasse の unit index に関する文献 (2012 年 1 月更新版)

[1] H. Hasse : Über die Klassenzahl abelscher Zahlkörper, Mathematische Lehrbucher und Monographien, Band I, Akademie-Verlag, Berlin, 1952, xii+190 S.; MR0049239 (14, 141a) (reviewed by O. F. G. Schilling; reprinted in Reviews in Number Theory, volume 5, R02-10) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0046.26003 (reviewed by S. Kuroda) [scanned image]; [PDF, GDZ]; [PDF], DigiZeitschriften; Nachdruck der ersten Aufgabe mit einer Einleitung von J. Martinet : Springer-Verlag, New York and Berlin, 1985, xxii+xii+190 S.; MR0842666 (87j:11122a) (reviewed by Editors) [html, PDF, DVI, PostScript, BibTeX]; Mathematische Lehrbucher und Monographien, II. Abteilung: Mathematische Monographien, Band 1, Akademie-Verlag, Berlin, 1985; MR0834791 (87j:11122b) (reviewed by Editors) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0668.12004 [html, PDF, XML, ASCII, DVI, PS, Bib-TeX], 0668.12003 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[2] K. Uchida : Imaginary abelian number fields with class number one, Tôhoku Math. J. (2) [東北數學雑誌第二輯] 24 (1972), no. 3, 487–499; doi:10.2748/tmj/1178241490, euclid.tmj/1178241490 [PDF (917 KB)]; 24_487.pdf (907 K), Journal@rchive, JST; MR0321904 (48 #269) (reviewed by T. Kubota; reprinted in Reviews in Number Theory, volume 5A, R208-205) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0248.12007 (Zusammenfassung des Autors) [scanned image].

[3] K. Yoshino and M. Hirabayashi : On the relative class number of imaginary abelian number field I, Mem. College Liberal Arts, Kanazawa Medical Univ. [金沢医科大学教養論文集] 9 (1981), 5–53; Zbl 0667.12003 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[4] K. Yoshino and M. Hirabayashi : On the relative class number of imaginary abelian number field II, Mem. College Liberal Arts, Kanazawa Medical Univ. [金沢医科大学教養論文集] 10 (1982), 33–81; Zbl 0667.12003 (reviewed by F. Halter-Koch) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[5] M. Hirabayashi : On the unit indices of imaginary abelian number fields I, Mem. Kanazawa Inst. Tech. [金沢工業大学研究紀要] 17 (1982), 1–4.

[6] M. Hirabayashi : On the unit indices of imaginary abelian number fields II, Mem. Kanazawa Inst. Tech. [金沢工業大学研究紀要] 17 (1982), 5–11.

[7] M. Hirabayashi : A Note on Unit Indices of CM-fields, Mem. Kanazawa Inst. Tech. [金沢工業大学研究紀要] 21 (1984), 173–176.

[8] M. Hirabayashi and K. Yoshino : The unit indices of imaginary abelian number fields, Proc. Japan Acad. Ser. A Math. Sci. [日本学士院紀要] 60 (1984), No. 6, 215–217; doi:10.3792/pjaa.60.215, euclid.pja/1195515067 [PDF (248 KB)]; MR0758046 (86i:11061) (from the introduction; reprinted in Reviews in Number Theory, volume 5B, R27-636) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0544.12002 [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[9] M. Hirabayashi : Remarks on Satz 29 of Hasse’s monograph, Mem. Kanazawa Inst. Tech. [金沢工業大学研究紀要] 26 (1987), 63–72.

[10] M. Hirabayashi and K. Yoshino : Remarks on unit indices of imaginary abelian number fields, Manuscripta Math. 60 (1988), fasc. 4, 423–436; DOI: 10.1007/BF01258662 [q19q72617951p385.pdf (440.5

KB)]; PPN365956996_0060_log30.pdf (925,99 KB),

GDZ; [PDF], DigiZeitschriften; MR0933473 (89e:11068) (reviewed by S. Shirai; reprinted in Reviews in Number Theory, volume 5B, R29-279) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0654.12002 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[11] M. Hirabayashi and K. Yoshino : Remarks on unit indices of imaginary abelian number fields II, Manuscripta Math. 64 (1989), fasc. 2, 235–251; DOI: 10.1007/BF01160122 [qk23168t071t4471.pdf (534.8

KB)]; PPN365956996_0064_log18.pdf (958,89 KB),

GDZ; [PDF], DigiZeitschriften; MR0998489 (90j:11115) (reviewed by H. M. Edger; reprinted in Reviews in Number Theory, volume 5B, R29-679) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0703.11054 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, Bib-TeX].

[12] M. Hirabayashi and K. Yoshino : Unit indices of imaginary abelian number fields of type (2, 2, 2), J. Number Theory 34 (1990), no. 3, 346–361; doi:10.1016/0022-314X(90)90141-D [PDF (587 K)]; MR1049510 (91e:11125) (reviewed by K. Nagata; reprinted in Reviews in Number Theory, volume 5B, R29-706) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0705.11065 (reviewed by K. Lakkis) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[13] S. Louboutin : On the class number one problem for non-normal quartic CM-fields, Tôhoku Math. J.

(2) [東北數學雑誌第二輯] 46 (1994), no. 1, 1– 12; doi:10.2748/tmj/1178225798, euclid.tmj/1178225798 [PDF (947 KB)]; 46_1.pdf (461 K), Journal@rchive, JST; MR1256724 (94m:11130) (reviewed by K. Yamamura; reprinted in Reviews in Number Theory, volume 5B, R29-894) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0796.11050 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[14] M. Hirabayashi : Unit indices of some imaginary composite quadratic fields, Pacific J. Math. 164 (1994), No. 1, 87–104; euclid.pjm/1102622282 [PDF (1164 KB), DjVu (253 KB)]; MR1267502 (95a:11097) (reviewed by S. Louboutin; reprinted in Reviews in Number Theory, volume 5B, R27-767) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0796.11049 (reviewed by M. Hirabayashi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[15] F. Lemmermeyer : Ideal class groups of cyclotomic fields. I, Acta Arith. LXXII (1995), fasc. 4, 347–359; aa7244.pdf, icm; MR1348202 (96h:11111) (reviewed by S. Louboutin; reprinted in Reviews in Number Theory, volume 5B, R18-692) [html, PDF, DVI, PostScript, Bib-TeX]; Zbl 0837.11059 (reviewed by T. Metsänkylä) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[16] M. Hirabayashi : Unit indices of some imaginary composite quadratic fields II, Pacific J. Math. 173 (1996), No. 1, 93–104; euclid.pjm/1102365844 [PDF (822 KB), DjVu (203 KB)]; MR1387792 (97d:11163) (reviewed by S. Louboutin) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0849.11084 (reviewed by M. Hirabayashi) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[17] S. Louboutin : Determination of all quaternion octic CM-fields with class number 2, J. London Math. Soc. (2) 54 (1996), Part 2, 227–238; doi:10.1112/jlms/54.2.227 [227.full.pdf]; MR1405052 (97g:11122) (reviewed by K. Yamamura) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0861.11064 (reviewed by S. Louboutin) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[18] S. Louboutin : Hasse unit indices of dihedral octic CM-fields, Math. Nachr. 215 (2000), Issue 1 , 107– 113; DOI: 10.1002/1522-2616(200007)215:1<107::AID-MANA107>3.0.CO;2-A [PDF (139 k)]; MR1768197 (2001f:11186) (reviewed by F. Lemmermeyer) [html, PDF, DVI, PostScript, BibTeX]; Zbl 0972.11105 (reviewed by F. Lemmermeyer) [html, PDF, XML, ASCII, DVI, PS, BibTeX].

[19] 中島匠一 : 有理数体の (2, 2) 拡大の類数, 第 7 回北陸数論研究集会報告集 [Proceedings of the 7th Workshop on Number Theory] (金沢大学サテライトプラザ, 2008) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2009, pp. 46–58.

[20] 山村健 : CM 体の Hasse の unit index に関する文献 (2008 年 10 月更新版), 第 7 回北陸数論研究集会報告集 [Proceedings of the 7th Workshop on Number Theory] (金沢大学サテライトプラザ, 2008) (平林幹人, 野村明人, 山下浩, 木村巌, 菅野孝史編), 2009, pp. 134–138.